# Pseudo-labelling meets Label Smoothing for Noisy Partial Label Learning

### v2

@@ Abstract
Partial label learning (PLL) is a weakly-supervised learning paradigm where each training instance is paired with a set of candidate labels (partial label), one of which is the true label. Noisy PLL (NPLL) relaxes this constraint by allowing some partial labels to not contain the true label, enhancing the practicality of the problem. Our work centres on NPLL and presents a minimalistic framework that initially assigns pseudo-labels to images by exploiting the noisy partial labels through a weighted nearest neighbour algorithm. These pseudo-label and image pairs are then used to train a deep neural network classifier with label smoothing. The classifier’s features and predictions are subsequently employed to refine and enhance the accuracy of pseudo-labels. We perform thorough experiments on seven datasets and compare against nine NPLL and PLL methods. We achieve state-of-the-art results in all studied settings from the prior literature, obtaining substantial gains in fine-grained classification and extreme noise scenarios. Further, we show the promising generalisation capability of our framework in realistic crowd-sourced datasets.
Keywords:
partial label learning weakly-supervised learning
@@ 1 Introduction
The effectiveness of contemporary deep learning methods heavily relies on the presence of high-quality annotated data, which is costly, arduous, labor-intensive, and time-consuming to gather. Weakly supervised learning offers a potential solution to this challenge by learning from partially labelled or noisily labelled examples. It has been widely studied in different forms, including multi-label learning [40], semi-supervised learning [27], noisy label learning [14] etc. Our paper focuses on a weakly supervised setting called Partial Label Learning (PLL), where each training instance is paired with a set of candidate labels (partial label), out of which one is the true label. PLL commonly occurs in crowd-sourced datasets, when the annotators are indicated to select several possible labels if they are uncertain about an instance. It naturally arises in several applications such as web-mining [9, 18], ecological informatics [2] and multi-modal learning [3, 21].
A fundamental limitation of PLL is the assumption that the correct label is always included in the partial label. To overcome this limitation and further the practicality, Noisy PLL (NPLL), also referred to as Unreliable Partial Label Learning (UPLL) [23] was proposed. NPLL allows some partial labels to not contain the correct label. Prior works [30, 23, 26, 15, 35] on NPLL have predominantly pursued a label disambiguation approach by maintaining and updating a distribution over label probabilities. To tackle noisy samples (samples whose correct label is not in the partial label), current methods apply a detection cum mitigation technique [30, 15, 26] which are known to cause error propagation due to the unavoidable detection errors [35]. Further, most state-of-the-art (SOTA) methods [30, 15, 35] require a two-stage training. The first stage involves warm-up training of the classifier, typically using a PLL algorithm which is followed by their proposed NPLL strategy. Determining the optimal duration for warm-up poses a challenge, since a too long or short warm-up period can negatively impact the performance [35]. Some methods also need a clean validation set to decide the warm-up duration [26, 15], which may not always be available.
@@
In a nutshell, our framework has a single trainable component (e.g. a ResNet-18 [8] model). It does away with complex multi-branch networks [30] and operates seamlessly without warm-up, resulting in faster training time and lower memory requirements. We perform comprehensive experiments on seven datasets in widely different settings (varying noise rates and number of candidate labels). Albeit minimalistic, PALS outperforms SOTA methods by a significant margin. In contrast to previous methods, PALS largely maintains its performance even as the noise rate increases. The performance improvements, compared to other methods, become more pronounced with an increase in both the number of classes and the proportion of noise in partial labels. In fine-grained classification, PALS yields significant improvements (outperforming all other methods by over 7pp). Finally, moving beyond the simulated PLL and NPLL benchmarks as in prior literature, we also show that PALS obtains favourable results on real world crowd sourced datasets [25]. In summary, our work makes the following contributions:
  • We propose PALS, a novel iterative pseudo-labelling based framework for NPLL which involves KNN based pseudo-labelling and label smoothing. To our knowledge, we are the first to showcase the effectiveness of label smoothing for NPLL.
  • We show notable gains over nine SOTA PLL and NPLL methods in both simulated benchmarks (as in prior literature) and real world crowd sourced datasets. We present thorough ablation studies and quantitative experiments to support our design choices and demonstrate the efficacy of our approach.
@@ 2.2 Deep Partial Label Learning
With the advent of Deep Learning, a variety of identification-based strategies that employ a neural network backbone have been proposed. [36] temporally ensemble the predictions at different epochs to disambiguate the partial label. [20] use self-training to update the model and progressively identify the true label. [7] formulate the generation process of partial labels and develop two classifiers: one, risk-consistent and the other, classifier-consistent. [32] propose a family of loss functions that generalize the different loss functions proposed by earlier works [11, 4, 20]. [29] present PiCO, a framework that learns discriminative representations by leveraging contrastive learning and prototypical classification. [33] present a simple idea that uses consistency regularization on the candidate set and supervised learning on the non-candidate set. [34] upgrade PiCO and introduce self-training and prototypical alignment to achieve noteworthy results. However, none of these methods account for the possibility of noise within partial labels, which is the primary focus of our work.
@@ 3.2.1 Pseudo-labelling
where $Z_{i}=\sum_{c}\hat{q}_{c}({\bm{x}}_{i})$ is the normalization factor. Using these posterior probabilities, we select a maximum of $m$ most reliable image-label pairs for each class where $m$ is the $\delta$-quantile of $[a_{1},a_{2},\ldots,a_{C}]$. We do this to ensure a near-uniform distribution of samples across all classes. $a_{1},a_{2},\ldots,a_{C}$ are defined as follows.
@@ 3.2.2 Noise robust learning
$\mathcal{L}_{ce}({\bm{x}}_{i},\mathbf{e}_{\tilde{y}_{i}}^{LS,r})=-(1-r)\log(h^{\tilde{y}_{i}}(f({\bm{x}}_{i})))-\frac{r}{C}\sum^{C}_{j=1}\log(h^{j}(f({\bm{x}}_{i})))$ (7)
@@ 3.2.3 Optional regularization components
$\mathcal{L}_{mix}({\bm{x}}_{i},{\bm{x}}_{j})=\alpha_{m}\mathcal{L}_{ce}({\bm{x}}_{ij}^{mix},\mathbf{e}_{\tilde{y}_{i}}^{LS,r})+(1-\alpha_{m})\mathcal{L}_{ce}({\bm{x}}_{ij}^{mix},\mathbf{e}_{\tilde{y}_{j}}^{LS,r})$ (9)
@@
$\mathcal{L}_{final}=\sum_{i,j,j^{\prime}\in[\tilde{n}]}\mathcal{L}_{mix}(aug_{w}({\bm{x}}_{i}),aug_{w}({\bm{x}}_{j}))+\mathcal{L}_{mix}(aug_{s}({\bm{x}}_{i}),aug_{s}({\bm{x}}_{j\prime}))$ (10)
@@ 3.2.4 Partial label augmentation
$Y^{t+1}_{i}=\begin{cases}Y_{i}\cup\{\operatorname*{arg\,max}_{c}\;h_{t}^{c}(f_{t}(aug_{w}({\bm{x}}_{i})))\},&\text{ if }\max_{c}h_{t}^{c}(f_{t}(aug_{w}({\bm{x}}_{i})))>\lambda^{t}\\ Y_{i},&\text{otherwise}.\par\end{cases}$ (11)
@@ 4.1 Experimental Setup
Datasets We construct the PLL and NPLL benchmark for four different datasets: CIFAR-10, CIFAR-100, CIFAR-100H [13] and CUB-200 [31]. We follow the same dataset generation process of [30]. The generation process is governed by two parameters: the partial rate ($q$) and the noise rate ($\eta$). The partial rate $q$ represents the probability of an incorrect label to be present in the candidate set and the noise rate $\eta$ represents the probability of excluding the correct label from the candidate set. For CIFAR10, we experiment with $q\in\{0.1,0.3,0.5\}$, and for other datasets (with a larger number of classes) we select $q$ from the set $\{0.01,0.03,0.05\}$. We conduct experiments using different values of $\eta\in\{0.1,0.3,0.4,0.5\}$, where values of $\eta$ exceeding 0.3 indicate high noise scenarios. Further, we show the effectiveness of our approach in real-world annotation settings by considering three crowdsourced datasets: Treeversity#6, Benthic, Plankton [25]. Treeversity is a publicly available dataset of plant images, crowdsourced by the Arnold Arboretum of Harvard University. Benthic is a collection of seafloor images containing flora and fauna. Plankton consists of underwater plankton images. Each image in these datasets has multiple annotations provided by domain experts.
@@
Table 1: Comparison of PALS with the previous state-of-the-art methods across various partial and noise rates.
@@
$\eta=0.1$ $\eta=0.3$ $\eta=0.1$ $\eta=0.3$ $\eta=0.1$ $\eta=0.3$
RC 80.87$\pm$0.30 75.24$\pm$0.17 79.69$\pm$0.37 71.01$\pm$0.54 72.46$\pm$1.51 49.74$\pm$0.70
LWS 82.97$\pm$0.24 74.28$\pm$0.79 80.93$\pm$0.28 69.70$\pm$0.72 70.41$\pm$2.68 39.42$\pm$3.09
FREDIS 90.57$\pm$0.23 84.35$\pm$0.20 89.02$\pm$0.15 81.02$\pm$0.60 87.42$\pm$0.21 65.15$\pm$0.13
PiCO 90.78$\pm$0.24 84.96$\pm$0.12 89.71$\pm$0.18 82.25$\pm$0.32 88.11$\pm$0.29 68.75$\pm$2.62
CRDPLL 93.48$\pm$0.17 86.19$\pm$0.48 92.73$\pm$0.19 83.40$\pm$0.14 91.10$\pm$0.07 73.78$\pm$0.55
IRNet 93.44$\pm$0.21 92.38$\pm$0.21 92.81$\pm$0.19 91.35$\pm$0.08 91.51$\pm$0.05 86.19$\pm$0.41
PiCO+ 94.48$\pm$0.02 94.43$\pm$0.19 94.02$\pm$0.03 92.94$\pm$0.24 93.56$\pm$0.08 88.21$\pm$0.37
UPLLRS 95.16$\pm$0.10 94.65$\pm$0.23 94.32$\pm$0.21 93.85$\pm$0.31 92.47$\pm$0.19 91.55$\pm$0.38
ALIM-Scale 95.71$\pm$0.01 95.35$\pm$0.13 95.31$\pm$0.16 94.36$\pm$0.03 94.71$\pm$0.04 90.63$\pm$0.10
ALIM-Onehot 95.83$\pm$0.13 95.75$\pm$0.19 95.52$\pm$0.15 94.67$\pm$0.21 95.19$\pm$0.24 92.26$\pm$0.29
PALS 96.28$\pm$0.05 95.90$\pm$0.20 95.96$\pm$0.08 95.43$\pm$0.17 95.52$\pm$0.13 94.18$\pm$0.10
@@
$\eta=0.1$ $\eta=0.3$ $\eta=0.1$ $\eta=0.3$ $\eta=0.1$ $\eta=0.3$
RC 52.73$\pm$1.05 45.77$\pm$0.31 52.15$\pm$0.19 43.92$\pm$0.37 46.62$\pm$0.34 40.31$\pm$0.55
LWS 56.05$\pm$0.20 45.71$\pm$0.45 53.59$\pm$0.45 42.20$\pm$0.49 45.46$\pm$0.44 33.60$\pm$0.64
FREDIS 64.73$\pm$0.28 59.42$\pm$0.18 67.38$\pm$0.40 60.86$\pm$0.72 66.43$\pm$0.22 56.15$\pm$0.18
PiCO 68.27$\pm$0.08 58.97$\pm$0.09 67.38$\pm$0.09 58.64$\pm$0.28 67.52$\pm$0.43 58.18$\pm$0.65
CRDPLL 68.12$\pm$0.13 62.94$\pm$0.28 67.53$\pm$0.07 61.79$\pm$0.11 67.17$\pm$0.04 61.03$\pm$0.43
IRNet 71.17$\pm$0.14 68.77$\pm$0.28 71.01$\pm$0.43 68.18$\pm$0.30 70.73$\pm$0.09 68.09$\pm$0.12
PiCO+ 75.04$\pm$0.18 71.79$\pm$0.17 74.68$\pm$0.19 69.97$\pm$0.01 73.06$\pm$0.16 67.56$\pm$0.17
UPLLRS 75.73$\pm$0.41 71.72$\pm$0.39 - - 74.73$\pm$0.24 70.31$\pm$0.22
ALIM-Scale 77.37$\pm$0.32 76.45$\pm$0.30 77.60$\pm$0.18 75.92$\pm$0.14 76.86$\pm$0.23 75.67$\pm$0.17
ALIM-Onehot 76.52$\pm$0.19 76.09$\pm$0.23 77.27$\pm$0.23 75.29$\pm$0.57 76.87$\pm$0.20 74.49$\pm$0.61
PALS 80.90$\pm$0.50 79.78$\pm$0.13 80.33$\pm$0.04 78.52$\pm$0.24 80.00$\pm$0.25 77.87$\pm$0.29
Table 2: PLL Experiments
@@
Table 3: Extreme noise experiments
@@
Table 4: Fine-grained datasets
@@
Table 5: Crowdsourced Datasets

### v3

@@ Abstract
We motivate weakly supervised learning as an effective learning paradigm for problems where curating perfectly annotated datasets is expensive and may require domain expertise such as fine-grained classification. We focus on Partial Label Learning (PLL), a weakly-supervised learning paradigm where each training instance is paired with a set of candidate labels (partial label), one of which is the true label. Noisy PLL (NPLL) relaxes this constraint by allowing some partial labels to not contain the true label, enhancing the practicality of the problem. Our work centres on NPLL and presents a framework that initially assigns pseudo-labels to images by exploiting the noisy partial labels through a weighted nearest neighbour algorithm. These pseudo-label and image pairs are then used to train a deep neural network classifier with label smoothing. The classifier’s features and predictions are subsequently employed to refine and enhance the accuracy of pseudo-labels. We perform thorough experiments on seven datasets and compare against nine NPLL and PLL methods. We achieve state-of-the-art results in all studied settings from the prior literature, obtaining substantial gains in the simulated fine-grained benchmarks. Further, we show the promising generalisation capability of our framework in realistic, fine-grained, crowd-sourced datasets.
@@ 1 Introduction
The effectiveness of contemporary deep learning methods heavily relies on the presence of high-quality annotated data. Although this might be available for typical image classification, curating labeled datasets for fine-grained classification problems is costly, arduous and requires expert knowledge. Weakly supervised learning offers a potential solution to this challenge by learning from partially labelled or noisily labelled examples. It has been widely studied in different forms, including multi-label learning [36], semi-supervised learning [23], noisy label learning [12] etc. Our paper focuses on a weakly supervised setting called Partial Label Learning (PLL), where each training instance is paired with a set of candidate labels (partial label), out of which one is the true label. A fundamental limitation of PLL is the assumption that the correct label is always included in the partial label. To overcome this limitation and further the practicality, Noisy PLL (NPLL), also referred to as Unreliable Partial Label Learning (UPLL) [19] was proposed. NPLL allows some partial labels to not contain the correct label. In fine-grained classification, annotators may provide multiple possible labels (e.g., multiple bird species that look similar), making NPLL techniques useful for learning from such ambiguous annotations [2].
Prior works [26, 19, 22, 13, 31] on NPLL have predominantly pursued a label disambiguation approach by maintaining and updating a distribution over label probabilities. To tackle noisy samples (samples whose correct label is not in the partial label), current methods apply a detection cum mitigation technique [26, 13, 22] which are known to cause error propagation due to the unavoidable detection errors [31]. Further, most state-of-the-art (SOTA) methods [26, 13, 31] require a two-stage training. The first stage involves warm-up training of the classifier, typically using a PLL algorithm which is followed by their proposed NPLL strategy. Determining the optimal duration for warm-up poses a challenge, since a too long or short warm-up period can negatively impact the performance [31]. Some methods also need a clean validation set to decide the warm-up duration [22, 13], which may not always be available.
@@
In a nutshell, our framework has a single trainable component (e.g. a ResNet-18 [7] model). It eliminates multi-branch networks [26] and operates seamlessly without warm-up, resulting in faster training time and lower memory requirements. We perform comprehensive experiments on seven datasets in widely different settings (varying noise rates and number of candidate labels). PALS outperforms the SOTA methods by a significant margin. In contrast to previous methods, PALS maintains its performance largely even as the noise rate increases. Performance improvements, compared to other methods, become more pronounced with increasing numbers of classes and the proportion of noise in partial labels.
In fine-grained classification, we evaluate PALS in both simulated and real-world crowd-sourced benchmarks. PALS achieves significant improvements in CUB-200, outperforming all other methods by over 5-16pp. Moving beyond the simulated PLL and NPLL benchmarks as in prior literature, we also show that PALS obtains consistent results on realistic, fine-grained, crowd-sourced datasets [21]. In line with prior works, PALS also benefits from standard regularization techniques such as Mix-up and Consistency regularization. However, PALS achieves better performance regardless of the presence or absence of regularization techniques. In summary, our work makes the following contributions:
  • We propose PALS, a novel iterative pseudo-labelling based framework for NPLL, which involves KNN based pseudo-labelling and label smoothing. To our knowledge, we are the first to showcase the effectiveness of label smoothing for NPLL.
  • We show notable gains over nine SOTA PLL and NPLL methods in the simulated benchmarks from prior literature. Further, we are the first to demonstrate the effectiveness of NPLL for fine-grained crowd-sourced datasets, with PALS consistently performing well across different datasets and annotator configurations.
  • We present thorough ablation studies and quantitative experiments to support our design choices and demonstrate the efficacy of our approach.
@@ 2.2 Deep Partial Label Learning
With the advent of Deep Learning, a variety of identification-based strategies that employ a neural network backbone have been proposed. [32] temporally ensemble the predictions at different epochs to disambiguate the partial label. [16] use self-training to update the model and progressively identify the true label. [6] formulate the generation process of partial labels and develop two classifiers: one, risk-consistent and the other, classifier-consistent. [28] propose a family of loss functions that generalize the different loss functions proposed by earlier works [9, 3, 16]. [25] present PiCO, a framework that learns discriminative representations by leveraging contrastive learning and prototypical classification. [29] present a technique that uses consistency regularization on the candidate set and supervised learning on the non-candidate set. [30] upgrade PiCO and introduce self-training and prototypical alignment to achieve noteworthy results. However, none of these methods account for the possibility of noise within partial labels, which is the primary focus of our work.
@@ 3.2.1 Pseudo-labelling
where $Z_{i}$ is the normalization factor such $\sum_{c}\hat{q}_{c}({\bm{x}}_{i})=1$. Using these posterior probabilities, we select a maximum of $m$ most reliable image-label pairs for each class where $m$ is the $\delta$-quantile of $[a_{1},a_{2},\ldots,a_{C}]$. We do this to ensure a near-uniform distribution of samples across all classes. $a_{1},a_{2},\ldots,a_{C}$ are defined as follows.
@@ 3.2.2 Noise robust learning
$\begin{split}\mathcal{L}_{ce}({\bm{x}}_{i},\mathbf{e}_{\tilde{y}_{i}}^{LS,r})=-(1-r)\log(h^{\tilde{y}_{i}}(f({\bm{x}}_{i})))\\ -\frac{r}{C}\sum^{C}_{j=1}\log(h^{j}(f({\bm{x}}_{i})))\end{split}$ (7)
@@ 3.2.3 Optional regularization components
$\begin{split}\mathcal{L}_{mix}({\bm{x}}_{i},{\bm{x}}_{j})=\alpha_{m}\mathcal{L}_{ce}({\bm{x}}_{ij}^{mix},\mathbf{e}_{\tilde{y}_{i}}^{LS,r})\\ +(1-\alpha_{m})\mathcal{L}_{ce}({\bm{x}}_{ij}^{mix},\mathbf{e}_{\tilde{y}_{j}}^{LS,r})\end{split}$ (9)
@@
$\begin{split}\mathcal{L}_{final}=\sum_{i,j,j^{\prime}\in[\tilde{n}]}\mathcal{L}_{mix}(aug_{w}({\bm{x}}_{i}),aug_{w}({\bm{x}}_{j}))\\ +\mathcal{L}_{mix}(aug_{s}({\bm{x}}_{i}),aug_{s}({\bm{x}}_{j\prime}))\qquad i,j,j^{\prime}\in[\tilde{n}]\end{split}$ (10)
@@ 3.2.4 Partial label augmentation
$Y^{t+1}_{i}=\begin{cases}\begin{split}Y_{i}\cup\{\operatorname*{arg\,max}_{c}\;h_{t}^{c}(f_{t}(aug_{w}({\bm{x}}_{i})))\},\\ \text{ if }\max_{c}h_{t}^{c}(f_{t}(aug_{w}({\bm{x}}_{i})))>\lambda^{t}\end{split}\\ Y_{i},\text{ otherwise}.\par\end{cases}$ (11)
@@ 4.1 Experimental Setup
Datasets We construct the PLL and NPLL benchmark for four different datasets: CIFAR-10, CIFAR-100, CIFAR-100H [11] and CUB-200 [27]. We follow the same dataset generation process of [26]. The generation process is governed by two parameters: the partial rate ($q$) and the noise rate ($\eta$). The partial rate $q$ represents the probability of an incorrect label to be present in the candidate set and the noise rate $\eta$ represents the probability of excluding the correct label from the candidate set. For CIFAR10, we experiment with $q\in\{0.1,0.3,0.5\}$, and for other datasets (with a larger number of classes) we select $q$ from the set $\{0.01,0.03,0.05\}$. We conduct experiments using different values of $\eta\in\{0.1,0.2,0.3,0.4,0.5\}$, where values of $\eta$ exceeding 0.3 indicate high noise scenarios. Further, we show the effectiveness of our approach in real-world annotation settings by considering three fine-grained crowdsourced datasets: Treeversity#6, Benthic, Plankton [21]. Treeversity is a publicly available dataset of plant images, crowdsourced by the Arnold Arboretum of Harvard University. Benthic is a collection of seafloor images containing flora and fauna. Plankton consists of underwater plankton images. Each image in these datasets has multiple annotations provided by domain experts.
@@
$\eta=0.1$ $\eta=0.2$ $\eta=0.3$ $\eta=0.1$ $\eta=0.2$ $\eta=0.3$ $\eta=0.1$ $\eta=0.2$ $\eta=0.3$
RC 80.87$\pm$0.30 78.22$\pm$0.23 75.24$\pm$0.17 79.69$\pm$0.37 75.69$\pm$0.63 71.01$\pm$0.54 72.46$\pm$1.51 59.72$\pm$0.42 49.74$\pm$0.70
LWS 82.97$\pm$0.24 79.46$\pm$0.09 74.28$\pm$0.79 80.93$\pm$0.28 76.07$\pm$0.38 69.70$\pm$0.72 70.41$\pm$2.68 58.26$\pm$0.28 39.42$\pm$3.09
FREDIS 90.57$\pm$0.23 88.01$\pm$ 0.19 84.35$\pm$0.20 89.02$\pm$0.15 86.14$\pm$0.17 81.02$\pm$0.60 87.42$\pm$0.21 80.81$\pm$0.99 65.15$\pm$0.13
PiCO 90.78$\pm$0.24 87.27$\pm$0.11 84.96$\pm$0.12 89.71$\pm$0.18 85.78$\pm$0.23 82.25$\pm$0.32 88.11$\pm$0.29 82.41$\pm$0.30 68.75$\pm$2.62
CRDPLL 93.48$\pm$0.17 89.13$\pm$0.39 86.19$\pm$0.48 92.73$\pm$0.19 86.96$\pm$0.21 83.40$\pm$0.14 91.10$\pm$0.07 82.30$\pm$0.46 73.78$\pm$0.55
IRNet 93.44$\pm$0.21 92.57$\pm$0.25 92.38$\pm$0.21 92.81$\pm$0.19 92.18$\pm$0.18 91.35$\pm$0.08 91.51$\pm$0.05 90.76$\pm$0.10 86.19$\pm$0.41
PiCO+ 94.48$\pm$0.02 94.74$\pm$0.13 94.43$\pm$0.19 94.02$\pm$0.03 94.03$\pm$0.01 92.94$\pm$0.24 93.56$\pm$0.08 92.65$\pm$0.26 88.21$\pm$0.37
UPLLRS 95.16$\pm$0.10 - 94.65$\pm$0.23 94.32$\pm$0.21 - 93.85$\pm$0.31 92.47$\pm$0.19 - 91.55$\pm$0.38
ALIM-Scale 95.71$\pm$0.01 95.50$\pm$0.08 95.35$\pm$0.13 95.31$\pm$0.16 94.77$\pm$0.07 94.36$\pm$0.03 94.71$\pm$0.04 93.82$\pm$0.13 90.63$\pm$0.10
ALIM-Onehot 95.83$\pm$0.13 95.86$\pm$0.15 95.75$\pm$0.19 95.52$\pm$0.15 95.41$\pm$0.13 94.67$\pm$0.21 95.19$\pm$0.24 93.89$\pm$0.21 92.26$\pm$0.29
PALS 96.28$\pm$0.05 96.17$\pm$0.18 95.90$\pm$0.20 95.96$\pm$0.08 95.76$\pm$0.12 95.43$\pm$0.17 95.52$\pm$0.13 95.89$\pm$0.14 94.18$\pm$0.10
@@
$\eta=0.1$ $\eta=0.2$ $\eta=0.3$ $\eta=0.1$ $\eta=0.2$ $\eta=0.3$ $\eta=0.1$ $\eta=0.2$ $\eta=0.3$
RC 52.73$\pm$1.05 48.59$\pm$1.04 45.77$\pm$0.31 52.15$\pm$0.19 48.25$\pm$0.38 43.92$\pm$0.37 46.62$\pm$0.34 45.46$\pm$0.21 40.31$\pm$0.55
LWS 56.05$\pm$0.20 50.66$\pm$0.59 45.71$\pm$0.45 53.59$\pm$0.45 48.28$\pm$0.44 42.20$\pm$0.49 45.46$\pm$0.44 39.63$\pm$0.80 33.60$\pm$0.64
FREDIS 64.73$\pm$0.28 64.53 $\pm$ 0.39 59.42$\pm$0.18 67.38$\pm$0.40 63.86$\pm$0.50 60.86$\pm$0.72 66.43$\pm$0.22 63.09 $\pm$ 0.22 56.15$\pm$0.18
PiCO 68.27$\pm$0.08 62.24$\pm$0.31 58.97$\pm$0.09 67.38$\pm$0.09 62.01$\pm$0.33 58.64$\pm$0.28 67.52$\pm$0.43 61.52$\pm$0.28 58.18$\pm$0.65
CRDPLL 68.12$\pm$0.13 65.32$\pm$0.34 62.94$\pm$0.28 67.53$\pm$0.07 64.29$\pm$0.27 61.79$\pm$0.11 67.17$\pm$0.04 64.11$\pm$0.42 61.03$\pm$0.43
IRNet 71.17$\pm$0.14 70.10$\pm$0.28 68.77$\pm$0.28 71.01$\pm$0.43 70.15$\pm$0.17 68.18$\pm$0.30 70.73$\pm$0.09 69.33$\pm$0.51 68.09$\pm$0.12
PiCO+ 75.04$\pm$0.18 74.31$\pm$0.02 71.79$\pm$0.17 74.68$\pm$0.19 73.65$\pm$0.23 69.97$\pm$0.01 73.06$\pm$0.16 71.37$\pm$0.16 67.56$\pm$0.17
UPLLRS 75.73$\pm$0.41 - 71.72$\pm$0.39 - - - 74.73$\pm$0.24 - 70.31$\pm$0.22
ALIM-Scale 77.37$\pm$0.32 76.81$\pm$0.05 76.45$\pm$0.30 77.60$\pm$0.18 76.63$\pm$0.19 75.92$\pm$0.14 76.86$\pm$0.23 76.44$\pm$0.12 75.67$\pm$0.17
ALIM-Onehot 76.52$\pm$0.19 76.55$\pm$0.24 76.09$\pm$0.23 77.27$\pm$0.23 76.29$\pm$0.41 75.29$\pm$0.57 76.87$\pm$0.20 75.23$\pm$0.42 74.49$\pm$0.61
PALS 80.90$\pm$0.50 80.45$\pm$0.49 79.78$\pm$0.13 80.33$\pm$0.04 79.40$\pm$0.50 78.52$\pm$0.24 80.00$\pm$0.25 79.08$\pm$0.22 77.87$\pm$0.29
Table 1: Comparison of PALS with the previous state-of-the-art methods across various partial and noise rates. ‘-’ indicates missing values due to the code being unavailable.
@@
Table 2: PLL Experiments. $\eta$ is set to 00.
@@
Table 3: Extreme noise experiments. $\eta$ values are above $0.3$
@@
Table 4: Simulated fine-grained datasets.